Sunland Corporation sells rock-climbing products and also operates an indoor climbing facility for climbing enthusiasts. During the last part of 2022, Sunland had the following transactions related to notes payable.

Sept. 1 Issued a $15,600 note to Pippen to purchase inventory. The 3-month note payable bears interest of 9% and is due December 1. (Sunland uses a perpetual inventory system.)
Sept. 30 Recorded accrued interest for the Pippen note.
Oct. 1 Issued a $19,200, 10%, 4-month note to Prime Bank to finance the purchase of a new climbing wall for advanced climbers. The note is due February 1.
Oct. 31 Recorded accrued interest for the Pippen note and the Prime Bank note.
Nov. 1 Issued a $24,000 note and paid $9,200 cash to purchase a vehicle to transport clients to nearby climbing sites as part of a new series of climbing classes. This note bears interest of 8% and matures in 12 months.
Nov. 30 Recorded accrued interest for the Pippen note, the Prime Bank note, and the vehicle note.
Dec. 1 Paid principal and interest on the Pippen note.
Dec. 31 Recorded accrued interest for the Prime Bank note and the vehicle note.


Prepare journal entries for the transactions noted above.
Answer #1
(a)
Date Account Titles
Sept. 1 Inventory $        15,600.00
        Notes Payable $            15,600.00
Sept.30 Interest Expense ($15,600 X .09 X 1/12) $             117.00
         Interest Payable $                 117.00
Oct. 1 Equipment $        19,200.00
         Notes Payable $            19,200.00
Oct. 31 Interest Expense [($19200 X .10 X 1/12)= 160 + $117] $             277.00
              Interest Payable $                 277.00
Nov. 1 Equipment $        33,200.00
           Notes Payable $            24,000.00
            Cash $              9,200.00
Nov. 30 Interest Expense   [($24000 X .08 X 1/12) = 160 + 160 + 117 $             437.00
              Interest Payable $                 437.00
Dec. 1 Notes Payable $        15,600.00
Interest Payable $             351.00
                      Cash $            15,951.00
Dec.31 Interest Expense ($160 + $160) 320
                     Interest payable

320